# Benchmarking Legal Reasoning on 340 Law Exams

### v3

@@ Abstract
Long-form legal reasoning remains a key challenge for large language models (LLMs) in spite of recent advances in test-time scaling. We introduce LEXam, a novel benchmark derived from 340 law exams spanning 116 law school courses across a range of subjects and degree levels. The dataset comprises 4,886 law exam questions in English and German, including 2,841 long-form, open-ended questions and 2,045 multiple-choice questions. Besides reference answers, the open questions are also accompanied by explicit guidance outlining the expected legal reasoning approach such as issue spotting, rule recall, or rule application. Our evaluation on both open-ended and multiple-choice questions present significant challenges for current LLMs; in particular, they notably struggle with open questions that require structured, multi-step legal reasoning. Moreover, our results underscore the effectiveness of the dataset in differentiating between models with varying capabilities. Adopting an LLM-as-a-Judge paradigm with rigorous human expert validation, we demonstrate how model-generated reasoning steps can be evaluated consistently and accurately. Our evaluation setup provides a scalable method to assess legal reasoning quality beyond simple accuracy metrics. Project page: https://lexam-benchmark.github.io/
@@ 1 Introduction
Legal reasoning is a critical frontier for large language models (LLMs) specifically and artificial intelligence (AI) at large, requiring specialized domain knowledge and advanced reasoning abilities such as precedent interpretation, statutory analysis, and legal inference. Despite progress in general reasoning, legal reasoning remains difficult and under-assessed in NLP research. Moreover, the legal domain is inherently high-stakes and a failure to thoroughly examine the capabilities and limitations of models could lead to serious real-world consequences [32, 43, 40]. As LLMs are used increasingly for various legal tasks, such as legal judgment prediction [79, 70, 29, 16], legal summarization [5, 4], legal interpretation [69, 18, 17, 38], legal case retrieval [39, 90, 41], legal argument classification [75, 10], and contract drafting [34], it is imperative to assess and understand their capabilities and limitations. A recent U.S. case, in which an attorney filed an LLM-generated brief citing fabricated cases [6], highlights the risks of undetected model errors and the need for robust evaluation.
@@
However, LLM benchmarking on deterministic tasks has so far mainly concentrated in STEM domains, such as math Olympiad problems [85, 37, 76, 66] and physics problems [22, 84, 11]. Reasoning models such as OpenAI o3 [57] and DeepSeek R1 [13] have demonstrated impressive performance in tasks dominated by deduction and established scientific rules. Nevertheless, their abilities in more nuanced and context-dependent domains, such as the law, remain poorly understood [19, 53]. Legal reasoning requires not only rigorous deductive and inductive logic but also its application to complex practical scenarios (often involving ill-defined problems), an area still under-explored in terms of LLM performance. In addition, STEM benchmarks can easily automate validation, for example, by checking the final numeric answers or passing a formal verifier. Following this outcome-based evaluation approach, previous legal reasoning benchmarks [25, 20, 10, 87] have primarily focused on evaluating the correctness of a model’s final outputs, without explicitly assessing the intermediate steps required to reach a legally correct answer. As a result, these benchmarks frequently overlook the underlying reasoning process, making it unclear why LLMs sometimes fail to arrive at the correct conclusion. This limitation can lead to tangible harm when such LLM-based systems are deployed in practical settings.
@@
Furthermore, while jurisdictions vary in authoritative sources and handling of precedent, they generally adhere to a common underlying approach in legal reasoning, though with significant differences between common law and continental law systems. Common law systems (found in countries like the United States, United Kingdom, and Australia) emphasize case precedent and judge-made law, while continental law systems (prevalent in most European countries, Latin America, and parts of Asia) rely primarily on comprehensive legal codes and statutes. Regardless of the difference between legal traditions, legal reasoning typically involves tasks such as identifying the legal issue based on given facts, determining the applicable legal rules, and applying these rules to the specific facts of the problem at hand. In addition, legal reasoning requires a rich variety of linguistic expressions that often vary depending on the specific legal systems involved. Previous studies [30, 68] show that LLMs exhibit cultural bias associated with the use of different languages. It is therefore essential to evaluate LLMs using multilingual legal data to better align with realistic needs.
@@
To address these gaps, we introduce LEXam, a multilingual legal reasoning benchmark designed to assess the process-based and outcome-based correctness of the answers generated by LLMs, as illustrated in Figure 1. Based on exams in Swiss, European, and international law, along with jurisprudence and legal history, LEXam provides a comprehensive framework for evaluating the legal reasoning capabilities of LLMs through problem sets available in both English and German. The benchmark comprises 2,841 long-form, open-ended questions and 2,045 multiple-choice questions (MCQs). Each open question is paired with reference answers and clear normative guidance, outlining the expected legal reasoning chain. This allows for thorough benchmarking of an LLM’s legal reasoning capabilities beyond merely evaluating final answers. Questions are sourced from 340 law exams across 116 courses at the Bachelor’s and Master’s levels taught by leading law professors at the University of Zurich Faculty of Law, one of Europe’s largest and most prestigious law schools (see Appendix B for further details). As a result, LEXam spans a wide range of legal areas both in domestic and international law, thus providing a robust benchmark for the evaluation of legal reasoning skills.
Experimental results show that state-of-the-art (SOTA) models struggle with multi-step reasoning and applying legal rules and principles to novel scenarios. LLMs particularly struggle with open questions requiring structured legal reasoning. Our LLM-as-a-Judge analysis shows that model-generated reasoning steps can be evaluated consistently and accurately, closely aligning with human expert assessments. This approach offers a scalable method for assessing legal reasoning quality beyond simple accuracy metrics. Our work provides both a resource for NLP research in the legal domain and a practical tool for educators evaluating AI capabilities.
@@
2 Related Work
2.1 International Benchmarks
Multiple international benchmarks for legal reasoning have been established, comprising various individual datasets, primarily sourced from China and the United States. Examples are LegalBench [26], LawBench [21], LexEval [35], LEXTREME [52] or LexGLUE [9]. While they propel the field forward, they still fall short in the complexity of the tasks evaluated. In contrast, we introduce very challenging law exam questions, often requiring detailed, long-form answers.
In Table 1, we compare LEXam to prior legal QA datasets. We demonstrate that LEXam a) covers a wide array of legal disciplines (78 subdomains), b) is the first general legal reasoning dataset containing challenging long-form ground-truth answers, and c) is the only dataset with an expert-tuned LLM-as-a-judge system for evaluation.
2.2 Swiss Benchmarks
Research on NLP benchmarks around Swiss legal data started with the task of legal judgment prediction, where models predict the case outcome based on facts [49, 50]. Subsequently, many other tasks have been studied on Swiss court rulings, such as explainability for judgment prediction [77], negation scope resolution [12], sentence boundary detection [7], citation extraction [64], case importance prediction [71], law area prediction [64], summarization of leading decisions [67], drafting of considerations [64], translation [53], and the analysis of gender distribution in court decisions [44]. Additionally, prior work studied the anonymization [51] and risk of re-identification of involved persons in court rulings [54].
Most prior tasks focus on content from court rulings and, to some extent, legislation. To our knowledge, no Swiss legal dataset from law school exams exists. In much prior work, labels have been extracted through regexes instead of manual human annotation [45]. While regexes allow for very large-scale datasets, the covered tasks may not be as natural compared to fully human-annotated datasets. In this work, we collect a large dataset of real-world university law school exams, including detailed reference responses written by leading law professors.
Housing Statutes [88] Unknown US English Statutory Housing Law 6,853 TF
Sara [28] Unknown US English Tax Liability 160 TF
Brazilian Bar Exams [15] Unknown Brazil Portuguese 17 law areas 2,130 MC
COLIEE [31] CC BY 4.0 Japan, Canada Japanese, English Japanese Civil Law, Canadian Case Law 1,774 MC
GLOBALCIT [78] CC BY 4.0 International English International Citizenship Law 9,310 MC
JecQA [89] CC BY-NC-ND China Chinese Chinese Law 21,072 MC
MMLU (Legal Subset) [27] CC BY 4.0 Primarily US English General Legal Knowledge 1,970 MC
Multistate Bar Exam [88] Unknown US English Bar Exam questions 1,195 MC
PrivacyQA [65] MIT Global English Privacy Law & Data Protection 1,750 LF (140)
@@ 3 Dataset
In Figure 1 we give a high-level overview of our dataset construction process, which we describe in more detail in the following sections.
@@ 3.1 Construction
The raw dataset is derived from 340 law exams that span 116 courses taught at the University of Zurich between 2016 and 2023 and was downloaded from their publicly accessible website. It contains a total of 3,628 questions, comprising 2,867 open questions, 398 MCQs, and 363 true/false questions (TFQs). Using the university’s own course descriptions for reference, three legal experts among the authors consolidated the original courses into 78 based on domain similarity. The use and publication of the dataset was approved by the dean of the law school.
@@ 3.1.1 Open Questions
We divided the open questions into test and development (dev) sets to support few-shot learning. Specifically, for each course containing at least ten open questions, we randomly selected five for the dev set and assigned the rest to the test set. This procedure yielded 300 questions from 60 courses in the dev set, and 2,541 in the test set. We provide a sample open question here:
{mtbenchbox}
Question:
A (domiciled in Switzerland), B (domiciled in Switzerland) and C (domiciled in France) enter into a contract. The contract contains an arbitration clause: “Any dispute […] shall be resolved either by arbitration in Switzerland or by French courts.” Is the PILA applicable to determine the validity of the arbitration agreement if B initiates arbitration against A?
{mtbenchbox}
Reference Answer (excerpt):
1. Art. 176(1) PILA:
… - Seat = “in Switzerland” → not specific → may lead to invalidity […]
2. Domicile: A and B both in CH → Chapter 12 PILA not applicable (SFT view) …
{mtbenchbox}
LLM Answer (excerpt):
… Since the seat is Switzerland, validity is governed by Swiss law (Art. 193 PILA).
Conclusion: PILA applies. Arbitration agreement valid under Swiss law.
Legal Provisions: - Art. 193 PILA …
{mtbenchbox}
LLM Judgment (excerpt):
Model incorrectly cites Art. 193 PILA (wrong chapter); should apply Art. 176.
Fails to assess whether Chapter 12 applies (all parties in CH), ignores SFT jurisprudence, and does not address ambiguity in clause.
Correctness Score: [[0.2]]
Abbreviated example for illustration purposes. The full version is provided in Appendix E.3.
@@ 3.1.2 Multiple-Choice Questions
The MCQs were manually parsed into question stems and statements by three authors, where the stem provides contextual information to evaluate the truth of each statement. For each stem, we randomly generate MCQs that contain two to five statements extracted from the original MCQ. Each question includes one correct answer and three distractors, randomly selected from the set of all possible incorrect answers. The set of incorrect answers comprises all answer combinations in which at least one statement is incorrect, according to the ground truth of each statement. This ensures a uniform number of answer choices per question and a baseline accuracy of approximately 25% for a random guesser.
To unify the evaluation metric across formats, we convert all TFQs into a multiple-choice format. First, each TFQ is manually parsed into a question stem (if applicable) and a statement. For questions that share the same stem, we group them and randomly generate MCQs with two to five associated statements. For stemless questions, we group them by domain and design templates to generate MCQs with two to five randomly selected statements. As with the original MCQs, each converted question includes four answer choices: one correct option and three randomly selected incorrect options from the entire pool of possible answers. This process yields a total of 1,660 MCQs. A sample MCQ can be found in Appendix E.5.
While evaluating LLMs using MCQs has become common practice due to the ease of interpreting results, previous research has demonstrated that LLMs can exhibit unstable and sometimes unpredictable behavior in response to changes in question formatting, phrasing, or the order of answer choices [80, 61, 2]. To further examine this issue, we constructed an auxiliary MCQ dataset comprising 385 questions, each containing five statements. We systematically vary the number of answer choices (4, 8, 16, and 32) while keeping both the question stem and the order of the statements constant. This perturbation setup aims to diagnose whether models select the correct answers based on genuine domain understanding or merely guess due to insufficiently challenging distractors.
@@ 3.2 Dataset Statistics and Structured Features
The LEXam dataset comprises 2,841 open questions, split into development (dev, 10.6%) and test (89.4%) sets, and 1,660 MCQs in a single set. Detailed per-course statistics and textual descriptions are in Section B.2. Figure 2 illustrates the data distributions.
@@ 4.1 Open Questions
Section E.1 exhibits the prompt to answer open questions. It first informs about the relevant knowledge domain with course name, and then provides standard guidance that law students follow when answering exam questions.
@@
Evaluating open questions in specialized domains like law is challenging. Lexical or shallow semantic metrics, such as BLEU [60], ROUGE [36], BERTScore [83], and AlginScore [82], may not accurately reflect answer quality, as high lexical or semantic overlap does not necessarily indicate correct legal reasoning, and vice versa. Legal reasoning often involves low lexical similarity [87]. LLM-as-a-Judge [86] may help, but whether a general-domain LLM is capable of judging the quality of legal responses remains an open question. To develop a more suitable LLM judge and verify its reliability, we proceed as follows: First, two of the authors, both with formal doctoral-level legal training, draft a specialized judging prompt, as presented in Section E.2. Second, they conduct a pilot study to iteratively optimize the prompt (with GPT-4o as the judging LLM) until they were satisfied with the judging performance. This pilot was conducted on a diverse sample of courses and a key challenge was to calibrate penalties, especially in cases where the LLM’s answer introduced plausible but wrong information not found in the reference answer. Through this process, several refinements were also made to the prompt. Third, we use the developed LLM judge to conduct a large-scale evaluation on all LLMs’ answers, and the experts conduct a large-scale post-experiment verification on the LLM judging quality (see Section 6).
@@ 4.2 Multiple-Choice Questions
Section E.4 shows the prompt used for answering MCQs. It first provides the course title to indicate the relevant domain, then guides the LLM through standard legal reasoning steps such as clarifying facts, identifying issues, and explaining rules.

### v4

# LEXam: Benchmarking Legal Reasoning on 340 Law Exams

@@ Abstract
Long-form legal reasoning remains a key challenge for large language models (LLMs) in spite of recent advances in test-time scaling. To address this, we introduce LEXam, a novel benchmark derived from 340 law exams spanning 116 law school courses across a range of subjects and degree levels. The dataset comprises 4,886 law exam questions in English and German, including 2,841 long-form, open-ended questions and 2,045 multiple-choice questions. Besides reference answers, the open questions are also accompanied by explicit guidance outlining the expected legal reasoning approach such as issue spotting, rule recall, or rule application. Our evaluation on both open-ended and multiple-choice questions present significant challenges for current LLMs; in particular, they notably struggle with open questions that require structured, multi-step legal reasoning. Moreover, our results underscore the effectiveness of the dataset in differentiating between models with varying capabilities. Deploying an ensemble LLM-as-a-Judge paradigm with rigorous human expert validation, we demonstrate how model-generated reasoning steps can be evaluated consistently and accurately, closely aligning with human expert assessments. Our evaluation setup provides a scalable method to assess legal reasoning quality beyond simple accuracy metrics.
We have open-sourced our code on GitHub and released our data on Hugging Face. Project page: https://lexam-benchmark.github.io/.
@@ 1 Introduction
Legal reasoning is a critical frontier for large language models (LLMs) specifically and artificial intelligence (AI) at large, requiring specialized domain knowledge and advanced reasoning abilities such as precedent interpretation, statutory analysis, and legal inference. Despite progress in general reasoning, legal reasoning remains difficult and under-assessed in NLP research. Moreover, the legal domain is inherently high-stakes and a failure to thoroughly examine the capabilities and limitations of models could lead to serious real-world consequences (Kleinberg et al., 2018; Medvedeva & Mcbride, 2023; Mahari et al., 2023). As LLMs are used increasingly for various legal tasks, such as legal judgment prediction (Wang et al., 2024a; Shui et al., 2023; Jiang & Yang, 2023; Deng et al., 2023), legal summarization (Bauer et al., 2023; Ash et al., 2024), legal interpretation (Savelka et al., 2023; Engel & McAdams, 2024; Dugac & Altwicker, 2025; Luo et al., 2025), legal case retrieval (Ma et al., 2024; Zhou et al., 2023; Mahari et al., 2024), legal argument classification (Thalken et al., 2023; Chlapanis et al., 2024), and contract drafting (Lam et al., 2023), it is imperative to assess and understand their capabilities and limitations.
@@
However, LLM benchmarking on deterministic tasks has so far mainly concentrated in STEM domains, such as math Olympiad problems (Zheng et al., 2022; Liu et al., 2023; Tsoukalas et al., 2024; Ren et al., 2025) and physics problems (Feng et al., 2025; Zhang et al., 2025; Chow et al., 2025). Reasoning models such as OpenAI o3 (OpenAI, 2025) and DeepSeek-R1 (DeepSeek-AI, 2025a) have demonstrated impressive performance in tasks dominated by deduction and established scientific rules. Nevertheless, their abilities in more nuanced and context-dependent domains, such as the law, remain poorly understood (Engel et al., 2025; Niklaus et al., 2025). Legal reasoning requires not only rigorous deductive and inductive logic but also their application to complex practical scenarios, often involving ill-defined problems. This domain offers an ideal testbed for non-formal reasoning, yet it remains underexplored in the context of LLM performance. In contrast, STEM benchmarks allow for straightforward validation—for example, by checking final numerical answers or employing formal verifiers. Adopting this outcome-based evaluation paradigm, prior legal reasoning benchmarks (Guha et al., 2023; Fei et al., 2023; Chlapanis et al., 2024; Zheng et al., 2025) have primarily assessed the correctness of a model’s final outputs, without explicitly evaluating the intermediate steps necessary for reaching a legally sound conclusion. Consequently, these benchmarks often overlook the reasoning process itself, leaving it unclear why LLMs fail to reach correct conclusions in some cases. This limitation raises concerns of tangible harm when LLM-based systems are deployed in practice.
@@
Furthermore, while jurisdictions differ in their authoritative sources and treatment of precedent, they share a common foundation in legal reasoning, though with notable distinctions between common law and continental law systems. Common law systems (e.g. the United States, United Kingdom, and Australia) emphasize case precedent and judge-made law, whereas continental law systems (prevalent in most of Europe, Latin America, and parts of Asia) rely primarily on comprehensive legal codes and statutes. Across traditions, legal reasoning involves identifying the legal issue from the facts, determining the relevant rules, and applying those rules to the case. It also depends on diverse linguistic expressions that vary across systems. Prior studies (Ryan et al., 2024; Jin et al., 2025; Fan et al., 2025) show that LLMs exhibit cultural biases tied to language use. Evaluating LLMs on multilingual legal data is therefore essential to align them with real-world needs.
@@
To address these gaps, we introduce LEXam, a multilingual legal reasoning benchmark designed to assess the process-based and outcome-based correctness of the answers generated by LLMs, as illustrated in Figure 1. Based on exams in Swiss, European, and international law, LEXam provides a comprehensive framework for evaluating the legal reasoning capabilities of LLMs through problem sets available in both English and German. The benchmark comprises 2,841 long-form, open-ended questions and 2,045 multiple-choice questions (MCQs). Each open question is paired with reference answers and clear normative guidance, outlining the expected legal reasoning chain. This allows for thorough benchmarking of an LLM’s legal reasoning capabilities beyond merely evaluating final answers. Questions are sourced from 340 law exams across 116 courses at the Bachelor’s and Master’s levels taught by leading law professors at the University of Zurich Faculty of Law, one of Europe’s largest and most prestigious law schools (see Appendix˜B for further details). As a result, LEXam spans a wide range of legal areas both in domestic and international law, thus providing a robust benchmark for the evaluation of legal reasoning skills.
Experimental results show that state-of-the-art (SOTA) models struggle with multi-step reasoning and applying legal rules and principles to novel scenarios. LLMs particularly struggle with open questions requiring structured legal reasoning. By designing and deploying an ensemble LLM-as-a-Judge evaluation framework, we demonstrate that model-generated reasoning steps can be evaluated consistently and accurately, closely aligning with human expert assessments. This approach provides a scalable method for assessing legal reasoning quality beyond simple accuracy metrics. Our work provides both a resource for NLP research in the legal domain and a practical tool for educators evaluating AI capabilities.
@@ 2 Dataset
Figure˜1 overviews our dataset construction process, detailed in the following sections.
@@ 2.1 Construction
The raw dataset is derived from 340 law exams that span 116 courses taught at the University of Zurich between 2016 and 2023 and was downloaded from their publicly accessible website. It contains a total of 3,628 questions, comprising 2,867 open questions, 398 MCQs, and 363 true/false questions (TFQs). Using the University’s own course descriptions for reference, three legal experts among the authors consolidated the original courses into 78 based on domain similarity. The use and publication of the dataset was approved by the dean of the law school. Figure˜2 provides a sample open question and a MCQ.
Figure 2: Illustration of a long-form open question (left, abbreviated example for illustration purposes. The full version is provided in Appendix E.3) and a MCQ with a set of candidate statements (right).
@@ 2.1.1 Open Questions
We divided the open questions into test and development (dev) sets to support few-shot learning. Specifically, for each course containing at least ten open questions, we randomly selected five for the dev set and assigned the rest to the test set. This procedure yielded 300 questions from 60 courses in the dev set, and 2,541 in the test set.
@@ 2.1.2 Multiple-Choice Questions
The MCQs were manually parsed into question stems and statements by three authors, where the stem provides contextual information to evaluate each statement. For each stem, we randomly generate MCQs that contain two to five statements extracted from the original MCQ. Each question includes one correct answer and three distractors, randomly selected from the set of all possible incorrect answers. The set of incorrect answers comprises all answer combinations where at least one statement is incorrect. This ensures a uniform number of answer choices per question and a baseline accuracy of approximately 25% for a random guesser.
To unify the evaluation metric across formats, we convert all TFQs into multiple-choice format. First, each TFQ is manually parsed into a question stem (if applicable) and a statement. For questions that share the same stem, we group them and randomly generate MCQs with two to five associated statements. For stemless questions, we group them by domain and design templates to generate MCQs with two to five randomly selected statements. As with the original MCQs, each converted question includes four answer choices: one correct option and three randomly selected incorrect options from the entire pool of possible answers. This process yields a total of 1,660 MCQs. A sample MCQ can be found in Figure 2.
While evaluating LLMs using MCQs has become common practice due to the ease of interpreting results, previous research has demonstrated that LLMs can exhibit unstable and sometimes unpredictable behavior in response to changes in question formatting, phrasing, or the order of answer choices (Wang et al., 2024b; Pezeshkpour & Hruschka, 2023; Alzahrani et al., 2024). To further examine this issue, we constructed an auxiliary MCQ dataset comprising 385 questions, each containing five statements. We systematically vary the number of answer choices (4, 8, 16, and 32) while keeping both the question stem and the order of the statements constant. This perturbation setup aims to diagnose whether models select the correct answers based on genuine domain understanding or merely guess due to insufficiently challenging distractors.
@@ 2.2 Dataset Statistics and Structured Features
The LEXam dataset comprises 2,841 open questions, split into dev (10.6%) and test (89.4%) sets, and 1,660 MCQs in a single set. Detailed per-course statistics and textual descriptions are in Section˜B.2. Figure 3 illustrates the data distributions.
@@ 3.1 Open Questions
Section˜E.1 exhibits the prompt to answer open questions. It first informs about the relevant knowledge domain with course name, and then provides standard guidance that law students follow when answering exam questions.
@@
Evaluating open questions in specialized domains like law is challenging. Lexical or shallow semantic metrics, such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), BERTScore (Zhang et al., 2020), and AlginScore (Zha et al., 2023), may not accurately reflect answer quality, as high lexical or semantic overlap does not necessarily indicate correct legal reasoning, and vice versa. Legal reasoning often involves low lexical similarity (Zheng et al., 2025). LLM-as-a-Judge (Zheng et al., 2023) may help, but whether a general-domain LLM is capable of judging the quality of legal responses remains an open question.
To develop a more suitable LLM judge and verify its reliability, we proceed as follows: First, two of the authors, both with formal doctoral-level legal training, draft a specialized judging prompt, as presented in Section˜E.2. Then, they conduct a pilot study to iteratively optimize the prompt (with GPT-4o as the judging LLM) until they were satisfied with the judging performance. This pilot was conducted on a diverse sample of courses and a key challenge was to calibrate penalties, especially in cases where the LLM’s answer introduced plausible but wrong information not found in the reference answer. Through this process, several refinements were also made to the prompt.
The next step is to identify capable LLMs for executing the expert-verified prompt. We rigorously evaluate various LLMs against human experts using the Alternative Annotator Test (Alt-test) (Calderon et al., 2025), detailed in Section 5. We find that only proprietary LLMs (e.g., GPT-4o, and Gemini-2.5-Pro) and extremely large reasoning models (e.g., DeepSeek-R1) consistently outperform human judges. Exclusive reliance on such models for evaluation, however, undermines the accessibility of LEXam. Encouragingly, we find that carefully constructed ensembles of open-source LLMs—for instance, taking the minimum score of Qwen3-32B and DeepSeek-V3—also surpass human judges. Our analyses further demonstrate that minimum-score ensembles of LLM judges consistently enhance evaluation performance. See Appendices G.4 and G.6 for intuition and detailed analyses of LLM Judge Ensembling. Appendices H and I report additional robustness checks for LLM judges.
In our evaluation, we adopt an minimum-score ensemble of GPT-4o, Qwen3-32B, and DeepSeek-V3 to grade open questions. Practitioners may instead rely on any LLM judge or ensemble that passes the Alt-test.
@@ 3.2 Multiple-Choice Questions
Appendix E.4 shows the prompt used for answering MCQs. It first provides the course title to indicate the relevant domain, then guides the LLM through standard legal reasoning steps such as clarifying facts, identifying issues, and explaining rules.